# GAEA: A Geolocation Aware Conversational Model

### v2

@@ 3.3 Question-Answer (QA) Pairs Generation
Figure 6: GAEA architecture with a single-stage training strategy including trainable MLP layers and LLM weights.
Reasoning Questions. We generate detailed image-caption QA pairs (Long-VQA) to enhance fine-grained reasoning in our GAEA model. We prompt GPT-4o [9] with each image, its scene labels, and country-specific geographical attributes, including GeoGuessr clues, traffic-side driving information, Köppen-Geiger climate zone, and land cover data. While scene labels are unique to each image, the other attributes provide country-level context. GPT-4o integrates this information to generate contextually rich and highly correlated captions with the provided geographic labels. These reasoning-based captions strengthen the model’s geolocalization and conversational capabilities and induce a rich semantic understanding in our model by infusing human-level cognition and inference capability, deducing the model to emphasize why particular image features might be associated with specific geographical contexts, reducing disinformation [43, 42]. In total, we curate 385K knowledge-driven LVQA pairs.
@@ 3.4 GAEA-Bench
Existing benchmarks for evaluating geolocalization tasks mainly focus on retrieval and classification-based methods, such as IM2GPS [24], IM2GPS3k [51], and GSW15k [18], which assess the distance between ground-truth and predicted GPS coordinates. However, there is a lack of conversational benchmarking datasets to evaluate the geolocalization and conversational capabilities of LMMs. We introduce GAEA-Bench, a geographically diverse and conversationally rich multimodal benchmark to address these shortcomings. GAEA-Bench is designed to assess LMMs across various question types, including MCQs, true/false, and long and short VQAs while integrating geolocalization tasks. It includes 4K image-text QA pairs that provide a rich geographical context for each image.
GAEA-Bench Curation. We curate a non-overlapping subset of highly geolocalizable MP-16 images, manually filtering out the non-geolocalizable ones. Using OpenStreetMaps (OSM), we generate metadata within a 1km radius and curate 1,000 short-form (SVQA), 1,000 multiple-choice (MCQ), and 1,000 true/false (T/F) questions. For long-form questions (LVQAs), we follow a similar process for generating reasoning questions in GAEA-Bench, resulting in an additional 1,000 questions. In total, we curate 4K diverse image-text QA pairs. To ensure that the GAEA-Bench remains independent of the training set, we select geographically distinct locations for its 4K samples. We show our GAEA-Bench annotation and curation process in Figure 3. The OSM metadata are fetched for each image and are passed to Qwen2.5-14B for generating several QA pairs, including SVQA, MCQ, and T/F.
4 GAEA Architecture
@@ 5 Benchmarking and Evaluations
Figure 7: Our classification accuracy pipeline evaluates city and country predictions by comparing them against ground truth annotations derived from GPS coordinates, with GPT-4o serving as the evaluator.
@@
GeoChat-7B [27] 24.06 ($\downarrow$ 59.71%) 16.38 ($\downarrow$ 67.96%) 54.00 ($\downarrow$ 27.90%) 32.20 ($\downarrow$ 58.98%) 31.66 ($\downarrow$ 52.07%)
LLaMA-3.2-Vision-11B [10] 47.45 ($\downarrow$ 20.55%) 26.84 ($\downarrow$ 47.50%) 52.40 ($\downarrow$ 30.04%) 47.20 ($\downarrow$ 39.87%) 43.47 ($\downarrow$ 34.20%)
LLaVA-Next-Mistral-7B [33] 47.57 ($\downarrow$ 20.34%) 23.20 ($\downarrow$ 54.63%) 28.90 ($\downarrow$ 61.42%) 56.70 ($\downarrow$ 27.77%) 39.09 ($\downarrow$ 40.83%)
GLM-4V-9B [22] 40.63 ($\downarrow$ 31.97%) 21.29 ($\downarrow$ 58.36%) 56.30 ($\downarrow$ 24.83%) 50.60 ($\downarrow$ 35.54%) 42.21 ($\downarrow$ 36.10%)
Phi-3.5-Vision-Instruct [8] 48.04 ($\downarrow$ 19.56%) 14.31 ($\downarrow$ 91.57%) 54.40 ($\downarrow$ 27.36%) 57.20 ($\downarrow$ 27.13%) 43.49 ($\downarrow$ 34.17%)
InternVL2-8B [57] 52.42 ($\downarrow$ 12.22%) 30.93 ($\downarrow$ 39.51%) 55.30 ($\downarrow$ 26.17%) 56.90 ($\downarrow$ 27.52%) 48.89 ($\downarrow$ 25.99%)
LLaVA-OV-7B [30] 53.34 ($\downarrow$ 10.68%) 29.32 ($\downarrow$ 42.66%) 57.60 ($\downarrow$ 23.10%) 56.10 ($\downarrow$ 28.54%) 49.09 ($\downarrow$ 25.69%)
Qwen2.5-VL [12] 54.84 ($\downarrow$ 8.17%) 35.14 ($\downarrow$ 31.27%) 47.20 ($\downarrow$ 36.98%) 59.10 ($\downarrow$ 24.71%) 49.07 ($\downarrow$ 25.72%)
Gemini-2.0-Flash [47] 55.47 ($\downarrow$ 7.12%) 34.72 ($\downarrow$ 32.09%) 56.20 ($\downarrow$ 24.97%) 56.10 ($\downarrow$ 28.54%) 50.62 ($\downarrow$ 23.37%)
GPT-4o-mini [9] 58.82 ($\downarrow$ 1.51%) 34.13 ($\downarrow$ 33.25%) 54.00 ($\downarrow$ 27.90%) 34.17 ($\downarrow$ 56.47%) 45.28 ($\downarrow$ 31.46%)
GPT-4o [9] 63.62 ($\uparrow$ 6.53%) 49.56 ($\downarrow$ 3.07%) 59.37 ($\downarrow$ 20.73%) 69.83 ($\downarrow$ 11.04%) 60.59 ($\downarrow$ 8.28%)
GAEA (Ours) 59.72 51.13 74.90 78.50 66.06
@@
Table 1: We benchmark 11 open-source and proprietary LMMs on GAEA-Bench. Notably, GAEA outperforms all open-source models and fares higher than the proprietary models on decision-making questions (MCQs and TFs). We provide the relative performance change for each model compared to GAEA.
GAEA-1.6M training set comprises four distinct question types: Multiple Choice Questions (MCQs), True/False (T/F), and Short and Long Visual Question Answering (VQA). The GAEA model is meticulously trained to ensure conversational fluency while possessing the capability to geolocalize visual samples. Current evaluation frameworks primarily focus on standard geo-localization datasets, measuring accuracy using distance-based metrics at various scales, including Street (1 km), City (25 km), Region (200 km), Country (750 km), and Continent (2,500 km). However, these methods fail to assess the conversational capabilities of LMMs. To address this gap, we define our evaluation process in three key dimensions: (a) Conversational accuracy, (b) Quantitative geo-localization accuracy, and (c) Classification accuracy.
5.1 Evaluation and Metrics
Conversational Evaluation. Most geolocation-specific models operate as “black box” systems, providing GPS coordinates without offering any reasoning or justification behind their outputs. In contrast, GAEA is the first model of its kind, explicitly trained on 1.6 million instructions, which include a significant number of knowledge-reasoning question-answer pairs. This enables GAEA to integrate world knowledge, such as geographical clues, conversational meta-tags, and advanced reasoning capabilities, making its geolocation predictions more transparent and insightful. To address the challenges of complex conversational evaluation, we benchmark 10 state-of-the-art open-source and closed-source LMMs on GAEA-Bench, which is meticulously curated to evaluate LMMs on diverse question types, including multiple-choice, true/false, and open-ended questions (short and long VQAs). See §Section 8.1 for the baselines used in this work.
We employ different prompts for each type of question. We use GPT-4o as a judge and prompt it to score responses to various types of questions with different criteria. We use accuracy for MCQs and T/F, correctness for SVQA, and consistency, relevance, and geographical correctness for long VQAs (LVQAs); see the evaluation pipeline in Figure 5. Here, correctness refers to how closely the model’s output matches the location and the correct answer in the ground-truth response [49]. For LVQA, the consistency metric evaluates the fluency and readability of the model’s prediction [44, 48, 49], while geographical correctness assesses whether the model’s prediction accurately identifies the correct city and country, directly matching the ground-truth answer. This is further discussed in §Section 8.2, and Figures 16, 20 and 21.
Classification Accuracy. Figure 7 illustrates the classification accuracy pipeline at the city and country levels. For this evaluation, we introduce three new datasets: GeoDE [41], DollarStreet [21], and CityGuessr68k [28]. From GeoDE, we sampled 22K images based on 16 meta-tags having geolocalizable features. From DollarStreet, we manually sampled 1.3K images, removing indoor and non-geolocalizable samples. Since its metadata contains only country-level information, we evaluate this dataset solely for country classification. Additionally, we use the validation set of 14K images from CityGuessr and all 22K GeoDE samples for city and country classification tasks.
5.2 Results and Discussion
GAEA-Bench Evaluation. Table 1 presents the per-model performance of 12 recent LMMs on GAEA-Bench. The results offer several insights: (i) Our proposed model, GAEA, achieves the highest average performance across decision-making questions (T/F and MCQs) and Short VQAs. Among proprietary models, GPT-4o [9] overall performs the best, with an accuracy of 60.59%, excelling particularly in Long VQAs—outperforming GAEA by 6.53% in this category. However, both open-source and proprietary models struggle with short-form questions. E.g., GPT-4o’s accuracy drops from 63.62% on long questions to 49.5% on short questions. (ii) GAEA outperforms all LMMs with an average accuracy of 66.06%, surpassing GPT-4o by 8.28% and outperforming the second-best open-source model, LLaVA-OneVision [30], by 25.69%. (iii) Several open-source models, including LLaMA-3.2-11B [20], GLM-4V-9B [22], and Phi-3.5-Vision [8], achieve comparable overall performance. (iv) LMMs perform better on decision-making questions (MCQs and T/F) than open-ended questions; see Figure 17. E.g., LLaVA-OneVision experiences a 57.8% drop in accuracy on SVQA compared to T/F questions. The low performance on free-form questions underscores the challenge of using short questions to effectively assess conversational capabilities in the GAEA-Bench. We provide qualitative comparisons with several LMMs in Figures 10, 14 and 15
@@ 5.2 Results and Discussion
Figure 8: Classification accuracy for both city and country labels, where GAEA establishes itself as a strong baseline, surpassing several recent LMMs in performance.
@@
Standard Geo-localization Evaluation Table 2 compares GAEA’s performance with various specialized encoder-only methods across three geolocalization benchmarks. While GAEA is trained on a large-scale conversational dataset with geolocalization capabilities, it achieves competitive results against specialized models. We also evaluate against GaGA [19], which is trained on a dataset five times larger than ours, on IM2GPS and IM2GPS3k. However, we exclude comparisons on GSW-15K due to differences in dataset curation. Following the guidelines of [18], we reconstruct the GSW-15K benchmark.
@@
[…]
GAEA (Ours) 41.4 54.4 74.3 87.8
@@
[…]
GAEA (Ours) 33.7 52.3 70.8 84.0
@@
In IM2GPS3k, GAEA achieves the second-highest scores after PIGEON, outperforming GaGA across all four distance thresholds. It surpasses GaGA by  2.5% at the 25km radius and 3.66% at the country level. Additionally, GAEA outperforms the specialized model GeoCLIP [50] across all thresholds, with a 1.5% higher score in the region category and  1% improvements at the city and country levels. In IM2GPS, GAEA outperforms GaGA at 25 km and 2,500 km, remains competitive at 200 km and 750 km, and slightly surpasses PIGEON at the city level while maintaining competitive performance across other thresholds. We also evaluate GAEA on GSW-15K, one of the most challenging datasets, which includes non-geolocalizable landmarks. GAEA outperforms GeoCLIP [50] and GeoDecoder [18] on city-level distance and achieves comparable performance at the region and country levels. Figure 8 presents GAEA’s Classification Accuracy on three new datasets: CityGuessr68k-val [28], GeoDE [41], and DollarStreet [21]. GAEA outperforms recent LMMs, including LLaVA-OneVision [30], InternVL [17], and GLM-4V-9B [22], on both city- and country-level classification. These results highlight GAEA’s extensive geographical coverage and strong geolocation capabilities.
@@ 6 Conclusion
We introduced GAEA, the first interactive conversational model with specialized geolocation capabilities, explicitly trained on a large-scale conversational dataset, GAEA-1.6M. We meticulously designed the dataset to enhance GAEA’s reasoning, conversational abilities, and geolocation accuracy. We curated geolocalizable images from MP-16, GLDv2, and CityGuessr68k, enriching them with auxiliary context and metadata, such as geographic clues, and climate zones. In addition to a high-quality instruction set, we present GAEA-Bench, a comprehensive benchmark that evaluates LMMs across multiple question types, including MCQs, True/False, short- and long-VQAs. Our results show that GAEA outperforms recent LMMs on GAEA-Bench, demonstrating strong geolocation and conversational capabilities by leveraging OpenStreetMap (OSM) data. These findings establish GAEA as a strong baseline for future research in geolocalization.

### v3

@@ 3.1.2 Question-Answer (QA) Pairs Generation
Reasoning Questions. We generate detailed image-caption QA pairs (Long-VQA) to enhance fine-grained reasoning in our GAEA model. We prompt GPT-4o [9] with each image, its scene labels, and country-specific geographical attributes, including GeoGuessr clues, traffic-side driving information, Köppen-Geiger climate zone, and land cover data. While scene labels are unique to each image, the other attributes provide country-level context. GPT-4o integrates this information to generate contextually rich and highly correlated captions with the provided geographic labels. These reasoning-based captions strengthen the model’s geolocalization and conversational capabilities and induce a rich semantic understanding in our model by infusing human-level cognition and inference capability, enabling the model to emphasize why particular image features might be associated with specific geographical contexts, reducing disinformation [45, 44]. In total, we curate 237k knowledge-driven LVQA pairs.
@@ 3.2 GAEA-Bench
Existing benchmarks for evaluating geolocalization tasks mainly focus on retrieval and classification-based methods, such as IM2GPS [26], IM2GPS3k [53], and GSW15k [18], which assess the distance between ground-truth and predicted GPS coordinates. However, there is a lack of conversational benchmarking datasets to evaluate the geolocalization and conversational capabilities of LMMs. We introduce GAEA-Bench, a geographically diverse and conversationally rich multimodal benchmark to address these shortcomings. GAEA-Bench is designed to assess LMMs across various question types, including MCQs, true/false, and long and short VQAs while integrating geolocalization tasks. It includes 3.5k image-text QA pairs that provide a rich geographical context for each image.
Figure 5: Overview of the GAEA model architecture and workflow. An input image is first processed by a Vision Transformer (ViT) encoder, whose output is projected through a visual projector to obtain visual embeddings. Simultaneously, the input text prompt is converted into text embeddings. The combined visual and textual embeddings are then fed into the Qwen2.5 LLM space, which generates a response based on the multimodal input. We follow the single-stage training approach, unfreezing MLP, and performing LoRA fine-tuning in the same stage.
GAEA-Bench Curation. We curate a non-overlapping subset of highly geolocalizable MP-16 images, manually filtering out the non-geolocalizable ones. Using OpenStreetMaps (OSM), we generate metadata within a 1km radius and curate 975 short-form (SVQA), 978 multiple-choice (MCQ), and 978 true/false (T/F) questions. For long-form questions (LVQAs), we follow a similar process for generating reasoning questions in GAEA-Bench, resulting in an additional 383 questions. In total, we curate 3,314 diverse image-text QA pairs. To ensure that the GAEA-Bench remains independent of the training set, we select geographically distinct locations for its 3.5k samples. We show our GAEA-Bench annotation and curation process in Figure 4. The OSM metadata are fetched for each image and are passed to Qwen2.5-14B for generating several QA pairs, including SVQA, MCQ, and T/F.
4 Model Architecture
@@ 5 Benchmarking and Evaluations
GAEA-1.4M training set comprises four distinct question types: Multiple Choice Questions (MCQs), True/False (T/F), and Short and Long Visual Question Answering (VQA). GAEA is meticulously trained to ensure conversational fluency while possessing the capability to geolocalize visual samples. Current evaluation frameworks primarily focus on standard geo-localization datasets, measuring accuracy using distance-based metrics at various scales, including Street (1 km), City (25 km), Region (200 km), Country (750 km), and Continent (2,500 km). However, these methods fail to assess the conversational capabilities of LMMs. To address this gap, we define our evaluation process in three key dimensions: (a) Conversational accuracy, (b) Quantitative geo-localization accuracy, and (c) Classification accuracy.
5.1 Evaluation and Metrics
Conversational Evaluation. Most geolocation-specific models operate as “black box” systems, providing GPS coordinates without offering any reasoning or justification behind their outputs. In contrast, GAEA is the first model of its kind, explicitly trained on 1.4 million instructions, which include a significant number of knowledge-reasoning question-answer pairs. This enables GAEA to integrate world knowledge, such as geographical clues, conversational meta-tags, and advanced reasoning capabilities, making its geolocation predictions more transparent and insightful. To address the challenges of complex conversational evaluation, we benchmark 12 state-of-the-art open-source and closed-source LMMs on GAEA-Bench, meticulously curated to evaluate LMMs on diverse question types, including multiple-choice, true/false, and open-ended questions (short and long VQAs). See §Section 9.1 for the baselines used in this work.
We employ different prompts for each type of question. We use GPT-4o as a judge and prompt it to score responses to various types of questions with different criteria. We use accuracy for MCQs and T/F, correctness for SVQA, and consistency, relevance, and geographical correctness for long VQAs (LVQAs); see the evaluation pipeline in Figure 6. Here, correctness refers to how closely the model’s output matches the location and the correct answer in the ground-truth response [51]. For LVQA, the consistency metric evaluates the fluency and readability of the model’s prediction [46, 50, 51], while geographical correctness assesses whether the model’s prediction accurately identifies the correct city and country, directly matching the ground-truth answer. This is further discussed in §Section 9.2, and Figure 12.
Figure 6: Evaluation pipeline for conversational benchmarking on GAEA-Bench, highlighting various question types we introduce in our GAEA-Bench. Each question type is evaluated with various defined criteria using GPT-4o as a judge. For instance, SVQA is evaluated against accuracy and correctness, and LVQA is evaluated on Consistency, Fluency, and Relevancy criteria.
Figure 7: Classification and distance threshold accuracy computation pipeline simultaneously evaluates geolocalization performance at city and country level by comparing model predictions with ground truth annotations derived from reverse-geocoding GPS coordinates and accuracy at different distance thresholds by geocoding predictions of the model.
Classification Accuracy. Figure 7 illustrates the classification accuracy at the city and country levels and distance threshold accuracy computation pipeline. For this evaluation, we benchmark on three datasets: GeoDE [43], DollarStreet [23], and CityGuessr68k [30]. From GeoDE, we sampled 22k images based on 16 meta-tags having geolocalizable features. From DollarStreet, we manually sampled 1.3k images, removing indoor and non-geolocalizable samples. Since its metadata contains only country-level information, we evaluate this dataset solely for country classification. Additionally, we use the validation set of 14k images from CityGuessr and all 22k GeoDE samples for city and country classification tasks.
5.2 Results and Discussion
GAEA-Bench Evaluation. Table 1 presents the performance of 12 recent LMMs on GAEA-Bench. The results offer several insights: (i) Our proposed model, GAEA, achieves the highest average performance across decision-making questions (T/F and MCQs) and Short VQAs. Among proprietary models, GPT-4o [9] overall performs the best, with an accuracy of 60.3%, excelling particularly in Long VQAs—outperforming GAEA by 1.9% in this category. However, both open-source and proprietary models struggle with short-form questions. E.g., GPT-4o’s accuracy drops from 66.3% on long questions to 49.5% on short questions. (ii) GAEA outperforms all LMMs with an average accuracy of 67.5%, surpassing GPT-4o by 7.2% and outperforming the second-best open-source model, LLaVA-OneVision [32], by 18.2%. (iii) Several open-source models, including LLaMA-3.2-11B [21], GLM-4V-9B [24], and Phi-3.5-Vision [8], achieve comparable overall performance. (iv) LMMs perform better on decision-making questions (MCQs and T/F) than open-ended questions; see Figure 17. E.g., LLaVA-OneVision experiences a 26.5% drop in accuracy on SVQA compared to MCQ questions. The low performance on free-form questions underscores the challenge of using short questions in effectively assessing conversational capabilities in the GAEA-Bench. Figures 14, 15 and 16 shows comparisons with several LMMs.
@@ 5.2 Results and Discussion
GeoChat-7B [29] 23.9-40.5 16.4-36.2 54.5-18.2 32.1-46.4 33.2-34.3
LLaVA-Next-Mistral-7B [35] 48.0-16.4 23.2-29.4 29.2-43.5 56.7-21.8 37.7-29.8
Phi-3.5-Vision-Instruct [8] 48.7-15.7 14.3-38.3 54.7-18.0 57.4-21.1 42.9-24.6
LLaMA-3.2-Vision-11B [10] 48.6-15.8 29.9-22.7 53.2-19.5 47.1-31.4 44.0-23.5
GLM-4V-9B [24] 41.6-22.8 29.7-22.9 56.7-16.0 50.6-27.9 45.2-22.3
Qwen2.5-VL [12] 57.0-7.4 29.9-22.7 48.1-24.6 59.4-19.1 47.1-20.4
InternVL2-8B [60] 54.5-9.9 31.0-21.6 55.7-17.0 56.8-21.7 48.6-18.9
LLaVA-OV-7B [32] 54.1-10.3 31.5-21.1 58.0-14.7 56.4-22.1 49.3-18.2
Gemini-2.0-Flash [49] 58.3-6.1 34.7-17.9 57.2-15.5 56.7-21.8 50.5-17.0
GPT-4o-mini [9] * 61.8-2.6 34.1-18.5 54.9-17.8 33.9-44.6 43.4-24.1
GPT-4o [9] * 66.3+1.9 49.5-3.1 59.4-13.3 69.6-8.9 60.3-7.2
GAEA (Ours) 64.4 52.6 72.7 78.5 67.5
@@
Table 1: We benchmark 11 open-source and proprietary LMMs on GAEA-Bench. Notably, GAEA outperforms all open-source models and fares higher than the proprietary models on decision-making questions (MCQs and TFs). We provide the relative performance change for each model compared to GAEA. *- We use GPT-4o as a judge for evaluation, and it has been documented that LLMs as judges prefer their long-form output [55, 61], hence the scores for these models are likely overestimated.
@@
Figure 8: Classification accuracy for both city and country labels, where GAEA surpasses several recent LMMs in performance.
@@
Standard Geo-localization Evaluation. Table 2 compares GAEA’s performance with various specialized encoder-only methods across three standard geolocalization benchmarks. While GAEA is trained on a large-scale conversational dataset with geolocalization capabilities, it achieves competitive results against specialized models. We evaluate against GaGA [20], which is trained on a dataset five times larger than ours, on IM2GPS and IM2GPS3k. However, we exclude comparisons on GWS15k due to differences in dataset curation. We contacted the authors of [18] for the original GWS15k benchmark for fair evaluation.
@@
[2pt/2.5pt] GeoReasoner [34] 24.9 48.1 65.8 82.3
GaGA [20] 38.8 54.8 75.1 87.7
GAEA (Ours) 43.0 57.4 77.2 89.5
@@
[2pt/2.5pt] GeoReasoner [34] 26.5 40.4 57.7 72.8
GaGA [20] 33.0 48.0 67.1 82.1
GAEA (Ours) 36.9 56.0 73.2 86.7
@@
On the IM2GPS3k benchmark, GAEA consistently outperforms all existing LMMs, including the domain-specialized PIGEON model [25], across all four distance thresholds. Notably, GAEA surpasses PIGEON by 2.2% at the 200 km threshold and by approximately 1.4% at the continent level. Compared to GeoCLIP [52], GAEA achieves gains of 5.3% at the region level and 2.5% at the city level. Additionally, GAEA significantly outperforms the open-source models, GaGA [20], with improvements of 6.1% at the country level and 8.0% at the regional level and GeoReasoner [34] with 10.5% at city, 15.6% at region, and 13.9% at a 2500 km threshold. On IM2GPS, GAEA outperforms both the specialized LMMs, GeoReasoner and GaGA on all four distance metrics. It also surpasses PIGEON and GeoCLIP models at the city level by  2.1% and  1.2%, respectively, while maintaining competitive performance across other thresholds. We also evaluate GAEA on GWS15k, one of the most challenging datasets, which includes non-geolocalizable landmarks. GAEA outperforms GeoCLIP [52] and GeoDecoder [18] on city-level distance and achieves comparable performance at the region and country levels. Figure 8 presents GAEA’s Classification Accuracy on three new datasets: CityGuessr68k-val [30], GeoDE [43], and DollarStreet [23]. GAEA outperforms recent LMMs, including LLaVA-OneVision [32], InternVL [17], and GLM-4V-9B [24], on both city- and country-level classification. These results highlight GAEA’s extensive geographical coverage and strong geolocation capabilities.
@@ 6 Conclusion
We introduced GAEA, the first interactive conversational model with specialized geolocation capabilities, explicitly trained on a large-scale conversational dataset, GAEA-1.4M. We meticulously designed the dataset to enhance GAEA’s reasoning, conversational abilities, and geolocation accuracy. We curated geolocalizable images from MP-16, GLDv2, and CityGuessr68k, enriching them with auxiliary context and metadata, such as geographic clues and climate zones. In addition to a high-quality instruction set, we present GAEA-Bench, a comprehensive benchmark that evaluates LMMs across multiple question types, including MCQs, True/False, short- and long-VQAs. Our results show that GAEA outperforms recent LMMs on GAEA-Bench, demonstrating strong geolocation and conversational capabilities by leveraging OpenStreetMap (OSM) data. These findings establish GAEA as a strong baseline for future research in geolocalization.
7 Acknowledgement